OpenStudy (anonymous):

Find the maximum and minimum values of the function f(x,y)= x^2 - 4xy + 2y^2 + 3x. I've calculated the first and second derivatives and found Hf(P) but i don't know what to do after that

OpenStudy (anonymous):

you have used partial derivatives right?

OpenStudy (amistre64):

the xy part tells me its a conic thats been rotated at the origin

OpenStudy (amistre64):

its elliptic if i see it right

OpenStudy (amistre64):

hyperbolic paraboloid .. i was close :)

OpenStudy (amistre64):

its only going to have a low spot i believe, the graph is a pretty much a piece of paper lifted at the corners

OpenStudy (amistre64):

find your partials at zero like you would any other "curve"

OpenStudy (amistre64):

the partial derivatives give you the equation for the normal at any given point

OpenStudy (anonymous):

\begin{eqnarray*} \frac{\partial f}{\partial x}&=&2x-4y+3 \\ \frac{\partial f}{\partial y}&=&-4x+4y \end{eqnarray*} Both of the above are zero at stationary values.

OpenStudy (amistre64):

I rule itself is foggy; fx(x,y) + fy(x,y) ...something fxy(x,y)

OpenStudy (amistre64):

Fxx*Fyy-(Fxy)^2 rules perhaps?

OpenStudy (anonymous):

he'll then need to find \[\frac{\partial^2f}{\partial x^2},\frac{\partial^2f}{\partial y^2}, \frac{\partial^2f}{\partial x\partial y}\] then evaluate \[\Delta = \Big(\frac{\partial^2f}{\partial x\partial y}\Big)^2-\Big(\frac{\partial^2f}{\partial x^2}\Big)\Big(\frac{\partial^2f}{\partial x^2}\Big)\] A point is maximum if \[\Delta<0 , \frac{\partial^2f}{\partial x^2}<0\] A point is minimum if \[\Delta<0 , \frac{\partial^2f}{\partial x^2}>0\] A point is saddle point if \[\Delta>0\]

OpenStudy (amistre64):

Fancy typing always makes it better :) those are them